# GnnX-Bench: Unravelling the Utility of Perturbation-based Gnn Explainers through In-depth Benchmarking

### v1

@@ 1 Introduction and Related Work
Gnns have shown state-of-the-art performance in various domains including social networks [32, 13], drug discovery [56, 35, 36], modeling of physical systems [44, 8, 9, 10], event detection [12, 25], and recommendation engines [58]. Unfortunately, like other deep-learning models, Gnns are black boxes due to lacking transparency and interpretability. This lack of interpretability is a significant barrier to their adoption in critical domains such as healthcare, finance, and law enforcement. In addition, the ability to explain predictions is critical towards understanding potential flaws in the model and generate insights for further refinement. To impart interpretability to Gnns, several algorithms to explain the inner workings of Gnns have been proposed. The diversified landscape of Gnn explainability research is visualized in Fig. 1. We summarize each of the categories below:
Gradient-based: The gradient-based explainers [33, 7] follow the idea of the rate of change being represented by gradients. Additionally, the gradient of the prediction with respect to the input represents the sensitivity of the prediction with respect to the input. This sensitivity gives the importance scores and helps in finding explanations.
In addition to the methodology employed by explanation algorithms, the type of explanation offered represents a crucial component. Explanations can be broadly classified into two categories: factual reasoning [59, 30, 40, 62, 18] and counterfactual reasoning [29, 43, 31, 6, 1, 50].
@@
Counterfactual explanations, on the other hand, elucidate why a particular prediction was not made by presenting alternative scenarios that could have resulted in a different decision. In the context of graphs, this involves identifying the smallest perturbation to the input graph that alters the prediction of the Gnn. Perturbations typically involves the removal of edges or modifications to node features. Counterfactual reasoners possess an additional advantage compared to factual reasoning, as they provide a recourse mechanism [46]. For instance, in the domain of drug discovery [21, 52], mutagenicity represents an undesirable property that impedes a molecule’s potential as a drug [23]. While factual explainers can attribute the subgraph responsible for mutagenicity, counterfactual reasoners can identify this subgraph along with the alterations required to render the molecule non-mutagenic.
1.1 Existing Benchmarking Studies on Gnn Explainability
GraphFrameX [4] and GraphXAI [2] represent two notable benchmarking studies. While both investigations have contributed valuable insights into Gnn explainers, certain unresolved investigative aspects persist.
Inclusion of counterfactual explainability: GraphFrameX and GraphXAI have focused on factual explainers for GNNs. [34] has discussed methods and challenges, but benchmarking on counterfactual explainers remains underexplored.
@@ 1.2 Contributions
In this benchmarking study, we systematically study perturbation-based factual and counter-factual explainers and identify their strengths and limitations in terms of their ability to provide accurate, meaningful, and actionable explanations for Gnn predictions. Overall, we make the following key contributions:
@@
Novel insights: The findings of our benchmarking study unveil stability to noise and variational factors, and generating feasible counterfactual recourses as two critical technical deficiencies that naturally lead us towards open research challenges.
Codebase: As a by-product, a meticulously curated, publicly accessible code base is provided (https://github.com/Armagaan/gnn-x-bench).
To keep the benchmarking study focused, we investigate only the perturbation-based explainability methods (highlighted in green in Fig. 1). While model-level explainers operate on the Gnn model, other forms of instance-level explainers yield diverse outputs spanning Directed Acyclic Graphs in PGMExplainer [47], model weights in Graph-Lime [18], node sets in Grad-CAM [33], among others. Consequently, enforcing a standardized set of inquiries across all explainers is not meaningful.
@@
| GNNExplainer [57] | Continuous relaxation | Mutual Information | Size | Yes | GC+NC | Transductive |
| PGExplainer [30] | Parameterized edge selection | Mutual Information | Size and/or connectivity | No | GC+NC | Inductive |
| TAGExplainer [51] | Sampling | Mutual Information | Size, Entropy | No | GC+NC | Inductive |
| GEM [27] | Granger Causality+Autoencoder | Causal Contribution | Size, Connectivity | No | GC+NC | Inductive |
| SubgraphX [62] | Monte Carlo Tree Search | Shapley Value | Size, connectivity | No | GC | Transductive |
| GstarX [63] | Monte Carlo sampling | HN-value | Size | No | GC | Inductive |
2 Preliminaries and Background
We use the notation to represent a graph, where denotes the set of nodes and denotes the set of edges. Each node is associated with a feature vector . We assume there exists a Gnn that has been trained on (or a set of graphs). Existing works has predominantly assumed the Gnn model to be a message-passing architecture such as Gcn [24], GraphSage [17], Gat [45], and Gin [53]. Therefore, we will base our subsequent discussion on this assumption.
@@ 2 Preliminaries and Background
where quantifies the distance between graphs and and is the set of all graphs one may construct by perturbing .
Typically, distance is measured in terms of number of edge perturbations while keeping the node set fixed. Under this assumption, and are defined as:
where denotes as adjacency matrix of .
@@ 2.1 Review of Perturbation-based Gnn Reasoning
Factual [61, 22]: The perturbation schema for factual reasoning usually consists of two crucial components: the subgraph extraction module and the scoring function module. Given an input graph , the subgraph extraction module extracts a subgraph ; and the scoring function module evaluates the model predictions
@@
For instance, GNNExplainer [57] identifies an explanation in the form of a subgraph that have the maximum influence on the prediction. In a follow-up work, PGExplainer [30] extends the same idea with an additional assumption of the graph to be a random Gilbert graph. Unlike the existing explainers, TAGExplainer [51] takes a two-step approach where the first step has an embedding explainer trained using a self-supervised training framework without any information of the downstream task.
A causality-based method GEM [27] uses the Granger causality to generate ground-truth explanations which are used to train the explainer.
SubgraphX [62] and GStarX [63] use cooperative game theoretic techniques. In particular, SubgraphX applies the Shapley value [41] to measure the importance of the subgraphs and GStarX uses HN values [16], to compute importance scores of a node for both graph and node classification tasks. Table 1 summarizes the key highlights.
Counterfactual [61, 22]: CF-GNNExplainer [29] aims to perturb the computational graph by using a binary mask matrix. The corresponding loss function quantifies the accuracy of the produced counterfactual, and captures the distance (or similarity) between the counterfactual graph and the original graph. In follow-up work, CF2 [43] extends this method by including a contrastive loss that jointly optimizes the quality of both the factual and the counterfactual explanation. RCExplainer [6], being both factual and counterfactual method, aims to identify a resilient subset of edges to remove such that it alters the prediction of the remaining graph. Finally CLEAR [31] generates counterfactual graphs by using a graph variational autoencoder. Table 2 summarizes the key highlights.
@@
| RCExplainer [6] | Instance level | GC+NC | Neural Network | Inductive |
| CF2 [43] | Instance level | GC+NC | Original graph | Transductive |
| CF-GNNExplainer [29] | Instance level | NC | Inference subgraph | Transductive |
| CLEAR [31] | Instance level | GC+NC | Variational Autoencoder | Inductive |
@@ 3 Benchmarking Framework
In this section, we outline the investigations we aim to conduct and the rationale behind them.
@@
We evaluate algorithms for both factual and counterfactual reasoning across a set of carefully chosen benchmark datasets (§ 4). Based on our holistic evaluation, we identify the pareto-optimal methods in terms of their performance, elucidating the trade-offs between different explainability techniques.
@@
Necessity and Reproducibility: The objective in this experiment is to quantify how central the explanation subgraph is for the Gnn towards making the prediction. We approach this question from the perspectives of necessity and reproducibility. Factual explanations are necessary if the removal of the explanation subgraph from the graph results in a significant decrease in prediction accuracy. Reproducibility [26], on the other hand, measures if the Gnn is retrained on the residual graph following the removal of the explanation, can it recover the original prediction?
@@
| Mutagenicity [37, 23] | 4337 | 131488 | 133447 | 14 | 2 | GC | F+CF |
| Proteins [11, 15] | 1113 | 43471 | 81044 | 32 | 2 | GC | F+CF |
| IMDB-B [55] | 1000 | 19773 | 96531 | 136 | 2 | GC | F+CF |
| AIDS [20] | 2000 | 31385 | 32390 | 42 | 2 | GC | F+CF |
| MUTAG [20] | 188 | 3371 | 3721 | 7 | 2 | GC | F+CF |
| NCI1 [48] | 4110 | 122747 | 132753 | 37 | 2 | GC | F |
| Graph-SST2 [61] | 70042 | 714325 | 644283 | 768 | 2 | GC | F |
| DD [15] | 1178 | 334925 | 843046 | 89 | 2 | GC | F |
| REDDIT-B [55] | 2000 | 859254 | 995508 | 3063 | 2 | GC | F |
| ogbg-molhiv [3] | 41127 | 1049163 | 2259376 | 9 | 2 | GC | CF |
| Tree-Cycles [57] | 1 | 871 | 1950 | 10 | 2 | NC | CF |
| Tree-Grid [57] | 1 | 1231 | 3410 | 10 | 2 | NC | CF |
| BA-Shapes [57] | 1 | 700 | 4100 | 10 | 4 | NC | CF |
@@ 4 Empirical Evaluation
Datasets: Table 3 showcases the principal statistical characteristics of each dataset employed in our experiments, along with the corresponding tasks evaluated on them. The Tree-Cycles, Tree-Grid, and BA-Shapes datasets serve as benchmark graph datasets for counterfactual analysis. These datasets incorporate ground-truth explanations [43, 27, 29]. Each dataset contains an undirected base graph to which predefined motifs are attached to random nodes, and additional edges are randomly added to the overall graph. The class label assigned to a node determines its membership in a motif. For more comprehensive information regarding the datasets, please refer to Appendix A.1.
Methods: The methods used in this benchmarking are delineated in the green branch at Figure 1.
Metrics: We use the following metrics in our study:
Size: In factual explanations, size denotes the number of edges in the explanation. In counterfactual, size denotes the number of edges perturbed to flip the label. Regardless of the explanation type, it is desirable for the explanation to be small.
Sufficiency (Fidelity): Sufficiency encodes the ratio of graphs for which the prediction derived from the explanation matches the prediction obtained from the complete graph [43]. Its value spans between 0 and 1. For factual explanations, higher values indicate superior performance, while in counter-factual lower is better since the objective is to flip the class label. Some works have used the term fidelity instead of sufficiency. In addition, some papers have reported necessity which is simply 1-sufficiency/fidelity.
Accuracy: Accuracy represents the percentage of correct explanations. Computing accuracy is feasible only when ground-truth explanations are available, which in our case restricts us to node classification in the three datasets of Tree-Cycles, Tree-Grid, and BA-Shapes. In line with the standards set by Cf, CF-GnnExplainer, and Gem, this metric pertains to the percentage of edges within the counterfactual that belong to the motif which decides the class.
@@ 4.1 Comparative Analysis
Across the majority of datasets, PGExplainer consistently delivers inferior results compared to other baseline methods. This behavior is more prominently visible in the spiderplot of the same data (See Fig. N in Appendix). However, it is challenging to identify a definitive best factual explainer, as no single technique dominates across all datasets. For instance, while RCExplainer performs exceptionally well in the Mutag dataset, it exhibits subpar performance in Imdb-B and Graph-SST2. Similar observations are also made for GnnExplainer in Reddit-B vs. Mutag and Nci1. Overall, we recommend using either RCExplainer or GNNExplainer as the preferred choices. The spider plot in Fig. N more prominently substantiates this suggestion.
@@
Counterfactual Explainers: We present separate analyses for explanations generated over graph classification and node classification tasks.
Table 4 presents the results on graph classification. RCExplainer is the best-performing explainer across majority of the datasets and metrics. However, it is important to acknowledge that RCExplainer’s sufficiency, when objectively evaluated, consistently remains high which is undesired. For instance, in the case of AIDS, the sufficiency of RCExplainer reaches a value of , signifying its inability to generate counterfactual explanations for of the graphs. This observation suggests that there exists considerable potential for further enhancement. We also note that while Clear achieves the best (lowest) sufficiency in AIDS, the number of perturbations it requires (size) is exorbitantly high to be useful in practical use-cases.
Table 5 presents the results on node classification. We observe that CF-GnnExplainer consistently outperforms Cf ( indicates the method to be entirely counterfactual). We note that our result contrasts with the reported results in Cf [43], where Cf was shown to outperform CF-GnnExplainer. Finally, when compared to graph classification, the sufficiency produced by the best methods in the node classification task is significantly lower indicating that it is an easier task. One possible reason might be the space of counterfactuals is smaller in node classification.
@@
| Method / Metric | Sufficiency | Size | Sufficiency | Size | Sufficiency | Size | Sufficiency | Size | Sufficiency | Size | Sufficiency | Size |
| RCExplainer () | 0.4 ±0.12 | 1.1 ±0.22 | 0.4 ±0.06 | 1.01 ±0.19 | 0.91 ±0.04 | 1.0 ±0.0 | 0.96 ±0.02 | 1.0 ±0.0 | 0.72 ±0.11 | 1.0 ±0.0 | 0.90 ±0.02 | 1 ±0.0 |
| CF | 0.90 ±0.12 | 1.0 ±0.0 | 0.50 ±0.05 | 2.78 ±0.98 | 0.98 ±0.02 | 5.25 ±0.35 | 1.0 ±0.0 | NA | 0.81 ±0.07 | 8.57 ±4.99 | 0.96 ±0.00 | 10.45 ±4.43 |
| Clear | 0.55 ±0.1 | 17.15 ±1.62 | OOM | OOM | 0.84 ±0.03 | 164.95 +- 47.93 | OOM | OOM | 0.96 ±0.02 | 218.62 ±0.0 | OOM | OOM |
@@
| Method / Metric | Sufficiency | Size | Acc.(%) | Sufficiency | Size | Acc.(%) | Sufficiency | Size | Acc.(%) |
CF-GnnExplainer |
0.5 ±0.08 | 1.03 ±0.16 | 100.0 ±0.0 | 0.09 ±0.06 | 1.42 ±0.55 | 92.70 ±4.99 | 0.37 ±0.05 | 1.37 ±0.59 | 91.5 ±4.36 |
Cf (=0) |
0.76 ±0.06 | 4.55 ±1.48 | 74.71 ±18.70 | 0.99 ±0.02 | 7.0 ±0.0 | 14.29 ±0.0 | 0.25 ±0.88 | 4.24 ±1.70 | 68.89 ±12.28 |

### v3

@@ 1 Introduction and Related Work
Gnns have shown state-of-the-art performance in various domains including social networks Manchanda et al. (2020); Chakraborty et al. (2023), biological sciences Ying et al. (2021); Rampášek et al. (2022); Ranjan et al. (2022), modeling of physical systems Thangamuthu et al. (2022); Bhattoo et al. (2022; 2023); Bishnoi et al. (2023), event detection Cao et al. (2021); Kosan et al. (2021) and traffic modeling Gupta et al. (2023); Jain et al. (2021); Wu et al. (2017); Li et al. (2020). Unfortunately, like other deep-learning models, Gnns are black boxes due to lacking transparency and interpretability. This lack of interpretability is a significant barrier to their adoption in critical domains such as healthcare, finance, and law enforcement. In addition, the ability to explain predictions is critical towards understanding potential flaws in the model and generate insights for further refinement. To impart interpretability to Gnns, several algorithms to explain the inner workings of Gnns have been proposed. The diversified landscape of Gnn explainability research is visualized in Fig. 1. We summarize each of the categories below:
Model-level: Model-level or global explanations are concerned with the overall behavior of the model and search for patterns in the set of predictions made by the model. XGNN Yuan et al. (2020), GLG-Explainer Azzolin et al. (2023), Xuanyuan et al. Xuanyuan et al. (2023), GCFExplainer Huang et al. (2023).
Instance-level: Instance-level or local explainers provide explanations for specific predictions made by a model. For instance, these explanations reason why a particular instance or input is classified or predicted in a certain way.
Gradient-based: They follow the idea of the rate of change being represented by gradients. Additionally, the gradient of the prediction with respect to the input represents the prediction sensitivity to the input. This sensitivity gives the importance scores and helps in finding explanations. SA and Guided-BP Baldassarre & Azizpour (2019), Grad-CAM Pope et al. (2019).
Decomposition-based: They consider the prediction of the model to be decomposed and distributed backward in a layer-by-layer fashion and the score of different parts of the input can be construed as its importance to the prediction. CAM and Excitation-BP Pope et al. (2019), GNN-LRP Schnake et al. (2021).
Perturbation-based: They utilize input perturbations to identify important subgraphs serving as factual or counterfactual explanations. GNNExplainer Ying et al. (2019b), PGExplainer Luo et al. (2020), SubgraphX Yuan et al. (2021), GEM Lin et al. (2021a), TAGExplainer Xie et al. (2022), CF Tan et al. (2022), RCExplainer Bajaj et al. (2021), CF-GNNexplainer Lucic et al. (2022), CLEAR Ma et al. (2022), Shan et al. (2021); Abrate & Bonchi (2021); Wellawatte et al. (2022)
Surrogate: They use the generic intuition that in a smaller range of input values, the relationship between input and output can be approximated by interpretable functions. The methods fit a simple and interpretable surrogate model in the locality of the prediction. GraphLime Huang et al. (2022), Relex Zhang et al. (2021), PGM-Explainer Vu & Thai (2020).
The type of explanation offered represents a crucial component. Explanations can be broadly classified into two categories: factual reasoning and counterfactual reasoning.
@@
Counterfactual explanations elucidate why a particular prediction was not made by presenting alternative scenarios that could have resulted in a different decision. In the context of graphs, this involves identifying the smallest perturbation to the input graph that alters the prediction of the Gnn. Perturbations typically involve the removal of edges or modifications to node features.
@@ 1.1 Contributions
In this benchmarking study, we systematically study perturbation-based factual and counterfactual explainers and identify their strengths and limitations in terms of their ability to provide accurate, meaningful, and actionable explanations for Gnn predictions. The proposed study surfaces new insights that have not been studied in existing benchmarking literature Amara et al. (2022); Agarwal et al. (2023)(See. App. J for details). Overall, we make the following key contributions:
@@
Novel insights: The findings of our benchmarking study unveil stability to noise and variational factors and generating feasible counterfactual recourses as two critical technical deficiencies that naturally lead us towards open research challenges.
Codebase: As a by-product, a meticulously curated, publicly accessible code base is provided (https://github.com/Armagaan/gnn-x-bench/).
2 Preliminaries and Background
We use the notation to represent a graph, where denotes the set of nodes and denotes the set of edges. Each node is associated with a feature vector . We assume there exists a Gnn that has been trained on (or a set of graphs).
@@ 2 Preliminaries and Background
| GNNExplainer | Continuous relaxation | Mutual Information | Size | Yes | GC+NC | Transductive |
| PGExplainer | Parameterized edge selection | Mutual Information | Size, Connectivity | No | GC+NC | Inductive |
| TAGExplainer | Sampling | Mutual Information | Size, Entropy | No | GC+NC | Inductive |
| GEM | Granger Causality+Autoencoder | Causal Contribution | Size, Connectivity | No | GC+NC | Inductive |
| SubgraphX | Monte Carlo Tree Search | Shapley Value | Size, Connectivity | No | GC | Transductive |
@@
where quantifies the distance between graphs and and is the set of all graphs one may construct by perturbing . Typically, distance is measured as the number of edge perturbations while keeping the node set fixed. In case of multi-class classification, if one wishes to switch to a target class label(s), then the optimization objective is modified as , where is the set of desired class labels.
@@ 2.1 Review of Perturbation-based Gnn Reasoning
Factual (Yuan et al. (2022); Kakkad et al. (2023)): The perturbation schema for factual reasoning usually consists of two crucial components: the subgraph extraction module and the scoring function module. Given an input graph , the subgraph extraction module extracts a subgraph ; and the scoring function module evaluates the model predictions
@@
For instance, while GNNExplainer Ying et al. (2019a) identifies an explanation in the form of a subgraph that have the maximum influence on the prediction, PGExplainer Luo et al. (2020) assumes the graph to be a random Gilbert graph. Unlike the existing explainers, TAGExplainer Xie et al. (2022) takes a two-step approach where the first step has an embedding explainer trained using a self-supervised training framework without any information of the downstream task. On the other hand, GEM Lin et al. (2021a) uses Granger causality and an autoencoder for the subgraph extraction strategy where as SubgraphX Yuan et al. (2021) employes a monte carlo tree search. The scoring function module uses mutual information for GNNExplainer, PGExplainer, and TAGExplainer. This module is different for GEM and SubgraphX, and uses casual contribution and Shapley value respectively. Table 1 summarizes the key highlights.
Counterfactual (Yuan et al. (2022)): The four major counterfactual methods are CF-GNNExplainer Lucic et al. (2022), CF2 Tan et al. (2022), RCExplainer Bajaj et al. (2021), and CLEAR Ma et al. (2022). They are instance-level explainers and apply to both graph and node classification tasks except for CF-GNNExplainer which is only applied to node classification. In terms of method, CF-GNNExplainer aims to perturb the computational graph by using a binary mask matrix. The corresponding loss function quantifies the accuracy of the produced counterfactual and captures the distance (or similarity) between the counterfactual graph and the original graph, whereas, CF2 Tan et al. (2022) extends this method by including a contrastive loss that jointly optimizes the quality of both the factual and the counterfactual explanation. Both of the above methods are transductive. As an inductive method, RCExplainer Bajaj et al. (2021), aims to identify a resilient subset of edges to remove such that it alters the prediction of the remaining graph while CLEAR Ma et al. (2022) generates counterfactual graphs by using a graph variational autoencoder. Table 2 summarizes the key highlights.
@@
| RCExplainer Bajaj et al. (2021) | Instance level | GC+NC | Neural Network | Inductive |
| CF2 Tan et al. (2022) | Instance level | GC+NC | Original graph | Transductive |
| CF-GNNExplainer Lucic et al. (2022) | Instance level | NC | Inference subgraph | Transductive |
| CLEAR Ma et al. (2022) | Instance level | GC+NC | Variational Autoencoder | Inductive |
| • : graph set. | |
| • : explanation subgraph of | |
| • : explanation set. | |
| • : residual graph set. | |
| • , , : the models trained on , , . All models are trained on the same labels. | |
| • : the prediction of the model on . | |
| • : the test accuracy of . |
@@ 3 Benchmarking Framework
In this section, we outline the investigations we aim to conduct and the rationale behind them. The mathematical formulation of the various metrics are summarized in Table 3.
@@
We evaluate algorithms for both factual and counterfactual reasoning and identify the pareto-optimal methods. The performance is quantified using explanation size and sufficiency Tan et al. (2022). Sufficiency encodes the ratio of graphs for which the prediction derived from the explanation matches the prediction obtained from the complete graph Tan et al. (2022). Its value spans between 0 and 1. For factual explanations, higher values indicate superior performance, while in counterfactual lower is better since the objective is to flip the class label.
@@
Necessity: Factual explanations are necessary if the removal of the explanation subgraph from the graph results in counterfactual graph (i.e., flipping the label).
Reproducibility: We measure two different aspects related to how central the explanation is towards retaining the prediction outcomes. Specifically, Reproducibility measures if the Gnn is retrained on the explanation graphs alone, can it still obtain the original predictions? On the other hand, Reproducibility measures if the Gnn is retrained on the residual graph constructed by removing the explanation from the original graph, can it still predict the class label? The mathematical quantification of these metrics is presented in Fig. 3.
@@
| Mutagenicity Riesen & Bunke (2008); Kazius et al. (2005) | 4337 | 131488 | 133447 | 14 | 2 | GC | F+CF |
| Proteins Borgwardt et al. (2005); Dobson & Doig (2003) | 1113 | 43471 | 81044 | 32 | 2 | GC | F+CF |
| IMDB-B Yanardag & Vishwanathan (2015) | 1000 | 19773 | 96531 | 136 | 2 | GC | F+CF |
| AIDS Ivanov et al. (2019) | 2000 | 31385 | 32390 | 42 | 2 | GC | F+CF |
| MUTAG Ivanov et al. (2019) | 188 | 3371 | 3721 | 7 | 2 | GC | F+CF |
| NCI1 Wale et al. (2008) | 4110 | 122747 | 132753 | 37 | 2 | GC | F |
| Graph-SST2 Yuan et al. (2022) | 70042 | 714325 | 644283 | 768 | 2 | GC | F |
| DD Dobson & Doig (2003) | 1178 | 334925 | 843046 | 89 | 2 | GC | F |
| REDDIT-B Yanardag & Vishwanathan (2015) | 2000 | 859254 | 995508 | 3063 | 2 | GC | F |
| ogbg-molhiv Allamanis et al. (2018) | 41127 | 1049163 | 2259376 | 9 | 2 | GC | CF |
| Tree-Cycles Ying et al. (2019a) | 1 | 871 | 1950 | 10 | 2 | NC | CF |
| Tree-Grid Ying et al. (2019a) | 1 | 1231 | 3410 | 10 | 2 | NC | CF |
| BA-Shapes Ying et al. (2019a) | 1 | 700 | 4100 | 10 | 4 | NC | CF |
@@ 4 Empirical Evaluation
Datasets: Table 4 showcases the principal statistical characteristics of each dataset employed in our experiments, along with the corresponding tasks evaluated on them. The Tree-Cycles, Tree-Grid, and BA-Shapes datasets serve as benchmark graph datasets for counterfactual analysis. These datasets incorporate ground-truth explanations Tan et al. (2022); Lin et al. (2021a); Lucic et al. (2022). Each dataset contains an undirected base graph to which predefined motifs are attached to random nodes, and additional edges are randomly added to the overall graph. The class label assigned to a node determines its membership in a motif.
@@ 4.1 Comparative Analysis
Insights: No single technique dominates across all datasets. For instance, while RCExplainer performs exceptionally well in the Mutag dataset, it exhibits subpar performance in Imdb-B and Graph-SST2. Similar observations are also made for GnnExplainer in Reddit-B vs. Mutag and Nci1. Overall, we recommend using either RCExplainer or GNNExplainer as the preferred choices. The spider plot in Fig. Q more prominently substantiates this suggestion. GNNExplainer is transductive, wherein it trains the parameters on the input graph itself. In contrast, inductive methods use pre-trained weights to explain the input. Consequently, transductive methods, such as GNNExplainer, at the expense of higher computation cost, has an inherent advantage in terms of optimizing sufficiency. Compared to other transductive methods, GNNExplainer utilizes a loss function that aims to increase sufficiency directly. This makes the method a better candidate for sufficiency compared to other inductive and transductive explainers. On the other hand, for RCExplainer, we believe that calculation of decision regions for classes helps to increase its generalizability as well as robustness.
@@
| Method / Metric | Suffic. | Size | Suffic. | Size | Suffic. | Size | Suffic. | Size | Suffic. | Size | Suffic. | Size |
| RCExplainer | ||||||||||||
| CF | NA | |||||||||||
| Clear | OOM | OOM | OOM | OOM | OOM | OOM | ||||||
Insights on graph classification (Table 5): RCExplainer is the best-performing explainer across the majority of the datasets and metrics. However, it is important to acknowledge that RCExplainer’s sufficiency, when objectively evaluated, consistently remains high, which is undesired. For instance, in the case of AIDS, the sufficiency of RCExplainer reaches a value of , signifying its inability to generate counterfactual explanations for of the graphs. This observation suggests that there exists considerable potential for further enhancement. We also note that while Clear achieves the best (lowest) sufficiency in AIDS, the number of perturbations it requires (size) is exorbitantly high to be useful in practical use-cases.
Insights on node classification (Table 6): We observe that CF-GnnExplainer consistently outperforms Cf ( indicates the method to be entirely counterfactual). We note that our result contrasts with the reported results in Cf Tan et al. (2022), where Cf was shown to outperform CF-GnnExplainer. A closer examination reveals that in Tan et al. (2022), the value of was set to , placing a higher emphasis on factual reasoning. It was expected that with , counterfactual reasoning would be enhanced. However, the results do not align with this hypothesis. We note that in Cf, the optimization function is a combination of explanation complexity and explanation strength. The contribution of is solely in the explanation strength component, based on its alignment with factual and counterfactual reasoning. The counterintuitive behavior observed with is attributed to the domination of explanation complexity in the objective function, thereby diminishing the intended impact of . Finally, when compared to graph classification, the sufficiency produced by the best methods in the node classification task is significantly lower indicating that it is an easier task. One possible reason might be the space of counterfactuals is smaller in node classification.
@@
| Method / Metric | Suffic. | Size | Acc.(%) | Suffic. | Size | Acc.(%) | Suffic. | Size | Acc.(%) |
CF-GnnEx |
Cf () |